OpenStudy (anonymous):

seven times some number added to 4 amounts to 31 added to the product of 4 and the number

OpenStudy (aum):

Let the number be x. "seven times some number added to 4 amounts to 31added to the product of 4 and the number" 7x + 4 = 31 + 4x solve for x subtract 4x 3x + 4 = 31 subtract 4: 3x = 27 divide by 3 x = 9 The number is 9.